How much will deposits of $35 made at the end of each month amount to after 10 years if interest is 6% compounded semi-annually?


1
Expert's answer
2021-12-08T11:33:06-0500

"FV=pmt.\frac{(1+i)^n-1}{i}"

"=35*\frac{(1+\frac{0.06}{6})^{10*12}-1}{\frac{0.06}{6}}"

=$8051.354